6x+2y=-10
2x+2y=-10

Answer by checkley71(8403)

6X+2Y=-10
-2X-2Y=10
-----------------------
4X=0
X=0 ANSWER.
6*0+2Y=-10
Y=-10/2
Y=-5 ANSWER.
-2*0-2*-5=10
10=10